The greatest value of the expression sin^2(15 π/8 -4x) – sin^2(17 π/8 -4x) for 0

Grade:11

1 Answers

Vikas TU
7 years ago
Greatest value for sin^2(15 π/8 -4x) will  be  in the interval x lying b/w 0 to pi/2.
.i,e 

15 π/8 -4x = pi/2
4x = 15pi/2 – pi/2
4x = 7pi
x = 7pi/4

and 
 
17 π/8 -4x = 0 {this should be minimum.)
4x = 17pi/8
x = 17pi/32